Grade 12Algebra

find the value of log3log3log9log729+3log3 basev3

Kumkum
12 Years agoGrade 12

1 Answer

Arun Kumar
ApprovedApproved Tutor Answer12 Years ago
Hello Student

9^3=729
log9(729)=3 3*logv3((v3)^2)=3*2=6
log3(3)=1
log3(1)=0
log3log3log9729+3logv33=6

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty